# Transfer learning for multifidelity simulation-based inference in cosmology

### v1

@@ Abstract
cosmology: cosmological parameters, large-scale structure of Universe, dark matter – methods: statistical
@@ 1 Introduction
Cosmological inference is increasingly turning to machine learning (ML) techniques for improved precision, accuracy, and efficiency. In particular, simulation-based inference (SBI) has emerged as a tool to enable statistical analysis of the large-scale structure beyond traditional Gaussian likelihood-based analysis. These techniques have been applied to weak lensing measurements of cosmic shear such as the Kilo-Degree Survey (KiDS, von Wietersheim-Kramsta et al., 2025; Lin et al., 2023), the Dark Energy Survey (DES, Jeffrey et al., 2021; Gatti et al., 2024; Jeffrey et al., 2025), and the Sloan Digital Sky Survey-III Baryon Oscillation Spectroscopic Survey (SDSS-III: BOSS, Lemos et al., 2024; Hahn et al., 2024).
@@
Our approach is straightforward: we perform transfer learning by first pre-training on a large corpus of cheaper, lower-fidelity data before training the model on a small set of accurate examples (a process known as fine-tuning). This widely used approach underpins foundation models, which are large, generic pre-trained models that can later be fine-tuned for specific tasks (He et al., 2016; Devlin et al., 2019; Dosovitskiy et al., 2021; Radford et al., 2021; Zhai et al., 2022; Kirillov et al., 2023). Pre-training allows models to learn generalisable features, which improves performance when adapting to new, related datasets (Bengio, 2012; Kornblith et al., 2019; Hoffmann et al., 2019; Mishra et al., 2022; Tahir et al., 2024; Lastufka et al., 2024). Some prior work has used this approach for cosmological inference (Sharma et al., 2024; Gondhalekar & Moriwaki, 2024), but there has been no comprehensive investigation into whether pre-training can substantially reduce the number of accurate simulations required to perform inference. Concurrent with this work, Krouglova et al. (2025) demonstrated that the exact same principle of transfer learning is effective for standard density estimation architectures such as neural spline flows, with applications directly to SBI.
This paper is structured as follows. Our methodology is described in Section 2. In Section 2.1 we describe the multifidelity simulation suites that we use for transfer learning. Section 2.2 presents the ML architectures and training procedures developed for this work, while the metrics used for model evaluation are introduced in Section 2.3. Section 3 presents the results of our multifidelity transfer learning methodology, and compares it with a high-fidelity-only approach for two examples: Section 3.1 explores a two-dimensional inference problem, and Section 3.2 addresses a more complicated five-dimensional inference problem, with a larger set of cosmological parameters and astrophysical nuisance parameters. Finally, we present a discussion of our results and our conclusions in Section 4.
@@ 2.1 Data
Figure 1: Comparison of the (scaled) dark matter density maps between paired simulations from the IllustrisTNG and $N$-body CMD simulation suites. Paired simulations have identical initial conditions as well as cosmological and astrophysical parameters. Panel a) shows a log-scaled dark matter density map, $M_{\text{cdm}}$, from the IllustrisTNG Latin hypercube (LH) suite. Panel b) gives the ratio between the IllustrisTNG $M_{\text{cdm}}$ and the paired $N$-body $M_{\text{cdm}}$, with a zoomed-in inlay of a particularly high contrast region. Panel c) shows the power spectrum ratio between the two maps, and panel d) shows the peak count statistics, computed by applying a four pixel $M_{\textrm{ap}}$ filter (see text for details) and computing the peak signal-to-noise ratio (SNR) over the background level. The exact cosmological and astrophysical feedback parameters are given in panel d).
@@
The CMD comprises thousands of simulations sampling universes with different cosmologies and astrophysical processes. These simulations are standardised to volumes of $(25\,h^{-1}\text{Mpc})^{3}$. For each simulation, 15 current-time ($z=0$) pseudo-independent 2D matter slices are extracted by considering 5 slices per dimension (of thickness $5\,h^{-1}\text{Mpc}$); these are then pixelised into bins of approximate area $(0.1\,h^{-1}\text{Mpc})^{2}$ to produce 2D images of the density fields with side of 256 pixels (see Villaescusa-Navarro et al., 2021 for the exact procedure). This work performs inference directly on density maps: we use 2D dark matter density maps, denoted as $M_{\text{cdm}}$, from the $N$-body simulations and the IllustrisTNG simulations. We then take the log of the maps and scale them such that the pixel statistics follow an approximate unit Gaussian distribution. These processed maps are then passed into the ML models. An example of these maps is given in Fig. 1, which shows the differences between two simulations, one hydrodynamical and one $N$-body, with identical parameters and initial conditions. Figure 1 explores the differences between the multifidelity simulations through a mass density ratio map, the power spectrum ratio, and a comparison of the peak statistics. The peak statistics were computed by first applying a four pixel aperture mass ($M_{\textrm{ap}}$) filter following Schneider et al. (1998), implemented using lenspack222 https://github.com/CosmoStat/lenspack.
@@
We apply our methodology to two of the simulation suites from the CMD: (i) the Latin hypercube suite (LH, Villaescusa-Navarro et al., 2022), which varies the matter density fraction $\Omega_{\textrm{m}}$ and the amplitude of the matter density power spectrum, parametrised by $\sigma_{8}$, alongside four astrophysical nuisance parameters ($A_{\textrm{SN,1}}$, $A_{\textrm{SN,2}}$, $A_{\textrm{AGN,1}}$, $A_{\textrm{AGN,2}}$); and, (ii) the Sobol28 suite (SB28, Ni et al., 2023), which varies $\Omega_{\textrm{m}}$, $\sigma_{8}$, the scalar spectral index $n_{\textrm{s}}$, the Hubble parameter $h$, and the baryonic density fraction $\Omega_{\textrm{b}}$, in addition to $23$ astrophysical nuisance parameters. The astrophysical parameters in the LH suite control the strength and behaviour of the stellar and active galactic nuclei (AGN) feedback in the simulations, while the SB28 suite varies a more detailed set of baryonic feedback processes controlling stellar and AGN feedback, supermassive black hole growth rates, star formation rates and stellar population modelling. It is worth noting that the small simulation box-size $25\,(h^{-1}\text{Mpc})^{3}$ may restrict the degree to which some of these effects, particularly AGN feedback, impact the simulations.
These suites contain 15000 and 30720 paired $N$-body and IllustrisTNG dark matter maps, respectively. In both cases we perform inference over only the cosmological parameters, implicitly marginalising over all nuisance parameters. For both suites, we use the last 200 cosmologies (corresponding to 3000 matter density maps) as holdout validation and test sets, ensuring no leakage between the training, validation, and test sets. We also augment the 2D matter map dataset by randomly flipping and rotating the images during training.
@@ 2.2 Model training
$=\mathbb{E}_{p(\theta,x)}\left[-\log q_{\varphi}(\theta\mid x)+\text{const.}\right].$. ] .
@@
The objective in Eq. 2 is used to first pre-train the network on $N$-body dark matter maps until convergence. It is then used again without modification to fine-tune the network on IllustrisTNG maps. We use this framework for simplicity, since only one network needs to be trained and all training can be done end-to-end. Extensions to e.g. neural likelihood estimation, and further variants, will be explored in future work. We do not envision these would require any significant modifications, but would need several training stages as in Jeffrey et al. (2021).
We performed an initial exploration of architectures for data compression and density estimation. We were partly motivated by the fact that common neural network architectures used for neural summarisation, such as CNNs, are well-suited for transfer between datasets. This stems from the inductive bias of CNNs, which encourages the learning of generic, transferable features that are then composed over a hierarchy of scales (Girshick et al., 2014; Yosinski et al., 2014; Kornblith et al., 2019). We found that the tailored CNN architecture from Villaescusa-Navarro et al. (2022) outperformed various standard architectures, such as ResNet (He et al., 2016) and ConvNext (Liu et al., 2022), both when pre-trained from natural image data or randomly initialised. We therefore used the CNN architecture from Villaescusa-Navarro et al. (2022) as our neural compression backbone, changing only the dimension of the final output layer to instead serve as a latent embedding. We found that using the CNN to compress matter density maps to larger latent dimension sizes slightly improved performance, so set the latent dimension to $128$.
We used a rational-quadratic neural spline flow (RQ-NSF, Durkan et al., 2019) as the NDE head of the network, implemented using the sbi Python package (Tejero-Cantero et al., 2022). We found this NDE architecture gave significant improvements over alternative popular choices, such as Masked Autoregressive Flows (MAFs, Papamakarios et al., 2017). We also found that inserting batch normalisation (Ioffe & Szegedy, 2015; Santurkar et al., 2018) layers between spline flow blocks (similar to the CNN architecture) substantially improved performance, both in pre-training and fine-tuning. An overview of the architecture is presented in Fig. 2.
@@
Once the architecture was selected, we performed end-to-end training of the network on the NPE objective in Eq. 2. We made a range of modifications that improved performance. We utilised a weight decay of $0.01$, which regularises the neural network by adding a small penalty term to the network weight magnitudes (Krogh & Hertz, 1991; Loshchilov & Hutter, 2017). This was particularly important for the fine-tuning stage, where regularisation while training on very small datasets was greatly beneficial. We used a short learning rate (LR) warm-up period, which has been found to improve deep learning model training (He et al., 2016; Goyal et al., 2017; Vaswani et al., 2017) with a number of posited explanations (see e.g., Gotmare et al., 2019; Kalra & Barkeshli, 2024). Larger batch-sizes (we selected 64) also improved performance for all models. We found that using a cyclic learning rate scheduler (Smith, 2017) as in Villaescusa-Navarro et al. (2022) improved performance when training on large datasets, i.e. $>\mathcal{O}(10^{5})$ maps. Baseline experiments and pre-training were performed with a LR of $2\times 10^{-4}$. Fine-tuning was performed with a LR of $1\times 10^{-5}$ and an exponential decay scheduler. All models were trained using the AdamW optimizer (Loshchilov & Hutter, 2017). All models pre-trained on $N$-body simulations used the entire training suite, and models with the lowest validation loss were saved for fine-tuning and evaluation.
@@ 2.3.1 Posterior Accuracy
Figure 3: Inference results on the IllustrisTNG LH suite for a 2D posterior over $\Omega_{\textrm{m}}$ and $\sigma_{8}$. We compare the performance of the two approaches: training with only high-fidelity maps (blue) against the transfer learning approach (green), which uses $13000$ $N$-body simulations for pre-training. An accurate benchmark model trained on the entire IllustrisTNG LH suite training set is shown by the dashed red line. Panel a) shows the mean test posterior probability (MTPP), panel b) shows the calibration error $\mathcal{C}$ of the modelled posterior, defined in Eq. 4, and panel c) shows the FoM, all as a function of high-fidelity dataset size $N$. Panel c) should be interpreted with the proviso that the model must be well-calibrated before the FoM measures genuine constraining power. All results show the mean and standard error over six independent training runs (including independent pre-training runs on the $N$-body simulations).
@@ 3 Results
We run a range of experiments comparing high-fidelity-only models against the transfer learning approach. We use $N$ to denote the number of IllustrisTNG 2D dark matter density maps used during the training stage. Of the dataset size $N$, $90\%$ is used as training data, while $10\%$ is used for validation data. We repeat all training runs six times, changing the initialisation of the network and the (random) order in which the training data is passed to the network. All results report the mean performance on the test set, using the lowest validation loss model from each repeat.
@@ 3.1 CAMELS Multifield Dataset: LH suite
We compare the performance of each method over a range of IllustrisTNG dataset sizes. We present the results of the LH experiments in Fig. 3. We find that small IllustrisTNG dataset sizes lead to poor performance when training from random network initialisation, whereas pre-training on $N$-body maps leads to good performance with very few simulations. For instance, a pre-trained model that is then fine-tuned with $N=200$ IllustrisTNG maps has higher MTPP than the high-fidelity-only approach with $N=6400$ maps.
@@
Insets: the overcoverage values per credibility level, see Eq. 4 and the surrounding text for details. The shaded orange region highlights the discrepancy between the ideal and observed distribution of credibility levels, which is quantified by the calibration error metric introduced in Eq. 4. Panel a) shows training from high-fidelity-only simulations, where models with small dataset sizes display significant overconfidence, and better calibration (though now mildly underconfident) is achieved at $N=6400$. Panel b) shows that transfer learning requires around $N=800$ IllustrisTNG maps to achieve good calibration.
@@
One complicating factor is the posterior calibration, quantified in Fig. 3b. This demonstrates that despite the high test posterior probability, models fine-tuned with very few IllustrisTNG maps appear to be poorly calibrated. We find that acceptable calibration is achieved after $N=800$ fine-tuning maps, while training with only high-fidelity maps requires $N=6400$ for similar posterior calibration. We therefore find at least a factor of $8$ reduction in the number of simulations required to produce a performant, well-calibrated model of the posterior.
We present calibration curves from a range of dataset sizes across the two approaches in Fig. 4. These show the standard cumulative distribution of observed credibility levels in the main panels, as well as the overcoverage distribution $\hat{p}_{i}/p_{i}$ in the insets. These reaffirm the calibration issues identified in Fig. 3: training from scratch with fewer than $N=6400$ maps leads to significantly overconfident posteriors. The high overcoverage at $\{0,1\}$ (paired with the below-ideal coverage in the middle of the distribution) is a clear indicator of overconfidence, since it indicates that the true parameters occur at extreme credibility levels too often. On the other hand, the transfer learning models display a more minor form of bias and overconfidence until reaching around $N=800$ maps.
The FoM performance as a function of dataset size is shown in Fig. 3c. When training using only high-fidelity simulations, low dataset sizes lead to low FoMs. This observation indicates that the features (or summary statistics) extracted by the CNN are not particularly informative. Additionally, the overconfident posteriors indicate limitations in the performance of the NDE, since it is incapable of producing trustworthy posteriors. We therefore conclude that both the CNN neural compressor and the density estimation model perform poorly with small training datasets.
@@
Three representative examples of inference on test cosmology maps are shown in Fig. 5. We compare models produced using only high-fidelity maps ($N=800$), transfer learning ($N=800$), and an “accurate” reference model trained on the full IllustrisTNG training set. These examples are qualitatively consistent with the analysis presented above. The high-fidelity-only $N=800$ model gives very uninformative constraints compared to the other two posteriors. On the other hand, the fine-tuned model appears well-calibrated, and only slightly less constraining than the model trained with ×$\times 16.25$ more high-fidelity maps. We present a similar comparison with a high-fidelity model trained on $N=3200$ in Appendix A, demonstrating that even for larger high-fidelity dataset sizes, pre-training yields significantly improved posteriors.
@@
Appendix B presents a range of further tests into the model performance. We found that the pre-trained models performed very poorly on high-fidelity maps when no fine-tuning was performed (corresponding to $N=0$). We explored the quality of the fine-tuned CNN compressor by freezing the CNN and re-training the NDE with larger dataset sizes. This presented more evidence that the limiting factor at very low fine-tuning dataset sizes was the NDE. We also showed that the small performance gap between the “accurate” model and the transfer learning models with larger dataset sizes was caused by slightly worse compression.
@@
Another consideration was the possibility that the paired aspect of the datasets was responsible for the significant performance gains from lower-fidelity pre-training. While we did make any explicit use of the simulation pairs, this could still have had an impact depending on the training dynamics of the network. We tested this in Appendix B and found strong evidence that the paired aspect of the source and target dataset has no impact on our results. We therefore conclude that our transfer learning approach does not depend on paired datasets. In principle, this means that large, pre-existing simulation suites could be used as multifidelity datasets without the need to pair initial conditions and cosmological parameters.
@@
We found that fine-tuning with the entire IllustrisTNG suite produced a slightly weaker model than the “accurate” benchmark model (i.e. the transfer learning curve in Fig. 3a does not intersect with the “accurate” performance). This is due to the low learning rate (LR) used during fine-tuning, and a higher LR recovered the “accurate” performance. This is consistent with the intuitive notion that a higher LR allows the training procedure to escape from the slightly sub-optimal region of the weight-space that is reached during pre-training. These observations are perfectly compatible with the study of Sharma et al. (2024), who found no clear benefits of transfer learning when using a large high-fidelity dataset.
@@
We present the degree of agreement between the posterior sample means $\hat{\theta}$ and the true cosmologies $\theta$ in Fig. 6, broken down for $\sigma_{8}$ and $\Omega_{\textrm{m}}$. These indicate pre-training yields very large improvements in the posterior for both parameters. $\Omega_{\textrm{m}}$ and $\sigma_{8}$ are inferred with similar precision relative to the baseline over the entire test set.

### v2

@@ Abstract
cosmology: cosmological parameters, large-scale structure of Universe, dark matter – methods: statistical - software: machine learning
@@ 1 Introduction
Cosmological inference is increasingly turning to machine learning (ML) techniques for improved precision, accuracy, and efficiency. In particular, simulation-based inference (SBI) has emerged as a tool to enable statistical analysis of the large-scale structure beyond traditional Gaussian likelihood-based analysis. These techniques have been applied to weak lensing measurements of cosmic shear such as the Kilo-Degree Survey (KiDS, von Wietersheim-Kramsta et al., 2025; Lin et al., 2023), the Dark Energy Survey (DES, Jeffrey et al., 2021; Gatti et al., 2024; Jeffrey et al., 2025), the Subaru Hyper Suprime-Cam (HSC, e.g. Novaes et al., 2025) and the Sloan Digital Sky Survey-III Baryon Oscillation Spectroscopic Survey (SDSS-III: BOSS, Lemos et al., 2024; Hahn et al., 2024; Thiele et al., 2024).
@@
Our approach is straightforward: we perform transfer learning by first pre-training on a large corpus of cheaper, lower-fidelity data before training the model on a small set of accurate examples (a process known as fine-tuning). This widely used approach underpins foundation models, which are large, generic pre-trained models that can later be fine-tuned for specific tasks (He et al., 2016; Devlin et al., 2019; Dosovitskiy et al., 2021; Radford et al., 2021; Zhai et al., 2022; Kirillov et al., 2023). Pre-training allows models to learn generalisable features, which improves performance when adapting to new, related datasets (Bengio, 2012; Kornblith et al., 2019; Hoffmann et al., 2019; Mishra et al., 2022; Tahir et al., 2024; Lastufka et al., 2024). Some prior work has used this approach for cosmological inference (Sharma et al., 2024; Gondhalekar & Moriwaki, 2024), but prior to this work there had been no comprehensive investigation into whether pre-training can substantially reduce the number of accurate simulations required to perform inference. Concurrent with this work, Krouglova et al. (2025) demonstrated that the exact same principle of transfer learning is effective for standard density estimation architectures such as neural spline flows, with applications directly to SBI. Since the submission of this manuscript, two independent studies proposed enhancements to a basic transfer learning strategy, each demonstrating improved outcomes (Thiele et al., 2025; Hikida et al., 2025).
This paper is structured as follows. Our methodology is described in Section˜2. In Section˜2.1 we describe the multifidelity simulation suites that we use for transfer learning. Section˜2.2 presents the ML architectures and training procedures developed for this work, while the metrics used for model evaluation are introduced in Section˜2.3. Section˜3 presents the results of our multifidelity transfer learning methodology, and compares it with a high-fidelity-only approach for two examples: Section˜3.1 explores a two-dimensional inference problem, and Section˜3.2 addresses a more complicated five-dimensional inference problem, with a larger set of cosmological parameters and astrophysical nuisance parameters. Finally, we present a discussion of our results and our conclusions in Section˜4.
@@ 2.1 Data
Figure 1: Comparison of the (scaled) dark matter density maps between paired simulations from the IllustrisTNG and NN-body CMD simulation suites. Paired simulations have identical initial conditions as well as cosmological and astrophysical parameters. Panel a) shows a log-scaled dark matter density map, $M_{\text{cdm}}$, from the IllustrisTNG Latin hypercube (LH) suite. Panel b) gives the ratio between the IllustrisTNG $M_{\text{cdm}}$ and the paired $N$-body $M_{\text{cdm}}$, with a zoomed-in inlay of a particularly high contrast region. Panel c) shows the power spectrum ratio between the two maps (blue line), along with a random selection of 250 spectral ratios between paired maps from the LH suite (grey lines). Panel d) shows the peak count statistics, computed by applying a four pixel $M_{\textrm{ap}}$ filter (see text for details) and computing the peak signal-to-noise ratio (SNR) over the background level. The exact cosmological and astrophysical feedback parameters are given in panel d).
@@
The CMD comprises thousands of simulations sampling universes with different cosmologies and astrophysical processes. These simulations are standardised to volumes of $(25\,h^{-1}\text{Mpc})^{3}$. For each simulation, 15 current-time ($z=0$) pseudo-independent 2D matter slices are extracted by considering 5 slices per dimension (of thickness $5\,h^{-1}\text{Mpc}$); these are then pixelised into bins of approximate area $(0.1\,h^{-1}\text{Mpc})^{2}$ to produce 2D images of the density fields with side of 256 pixels (see Villaescusa-Navarro et al., 2021 for the exact procedure). This work performs inference directly on density maps: we use 2D dark matter density maps, denoted as $M_{\text{cdm}}$, from the $N$-body simulations and the IllustrisTNG simulations (though we demonstrate our conclusions are unchanged for total matter density maps $M_{\text{tot}}$ in Appendix˜B). We then take the log of the maps and normalise them by subtracting the mean and dividing by the standard deviation of the pixel values, so that the resulting distribution is approximately unit Gaussian. These processed maps are then passed into the ML models. An example of these maps is given in Fig.˜1, which shows the differences between two simulations, one hydrodynamical and one $N$-body, with identical parameters and initial conditions. Figure˜1 explores the differences between the multifidelity simulations through a mass density ratio map, the power spectrum ratio, and a comparison of the peak statistics. The peak statistics were computed by first applying a four pixel aperture mass ($M_{\textrm{ap}}$) filter following Schneider et al. (1998), implemented using lenspack [^2] https://github.com/CosmoStat/lenspack.
@@
We apply our methodology to two of the simulation suites from the CMD: (i) the Latin hypercube suite (LH, Villaescusa-Navarro et al., 2022), which varies the matter density fraction $\Omega_{\textrm{m}}$ and the amplitude of the matter density power spectrum, parametrised by $\sigma_{8}$, alongside four astrophysical nuisance parameters ($A_{\textrm{SN,1}}$, $_{\textrm{SN,2}}, AA$$_{\textrm{AGN,1}}$, $A_{\textrm{AGN,2}}$); and, (ii) the Sobol28 suite (SB28, Ni et al., 2023), which varies $\Omega_{\textrm{m}}$, $\sigma_{8}$, the scalar spectral index $n_{\textrm{s}}$, the Hubble parameter $h$, and the baryonic density fraction $\Omega_{\textrm{b}}$, in addition to $23$ astrophysical nuisance parameters. The astrophysical parameters in the LH suite control the strength and behaviour of the stellar and active galactic nuclei (AGN) feedback in the simulations, while the SB28 suite varies a more detailed set of baryonic feedback processes controlling stellar and AGN feedback, supermassive black hole growth rates, star formation rates and stellar population modelling. It is worth noting that the small simulation box-size $25\,(h^{-1}\text{Mpc})^{3}$ may restrict the degree to which some of these effects, particularly AGN feedback, impact the simulations.
These suites contain 15000 and 30720 paired $N$-body and IllustrisTNG dark matter maps, respectively. In both cases we perform inference over only the cosmological parameters, implicitly marginalising over all nuisance parameters. For each suite, we reserve the last 200 cosmologies (corresponding to 3,000 matter density maps) as a holdout set, split evenly into 100 cosmologies (1,500 maps) for validation and 100 cosmologies (1,500 maps) for testing. The validation set is used for hyperparameter tuning and selection of the best model during training; the test set is used exclusively for model evaluation, with results reported in Section˜3. This strict split ensures no data leakage between the training, validation, and test sets. The paired nature of the CMD means our transfer learning approach inherently uses seed-matched simulations at both low and high fidelity in the training set. However, we explicitly verify that we achieve the same results with unpaired low and high fidelity datasets in Appendix˜B.
We augment the 2D matter map dataset by randomly flipping and rotating the images during training. We perform no further processing or augmentation, such as field smoothing or addition of noise, to the low or high fidelity maps.
@@ 2.2 Model training
$=\mathbb{E}_{p(\theta,x)}\left[-\log q_{\varphi}(\theta\mid x)+\text{const.}\right].$
@@
The objective in Eq.˜2 is used to first pre-train the network on $N$-body dark matter maps until convergence. It is then used again without modification to fine-tune the network on IllustrisTNG maps. All weights from the pre-trained network are transferred and used to initialise the fine-tuning stage on the high-fidelity maps, with the entire network being trained during this phase. We do not freeze any layers or restrict fine-tuning to specific components (e.g. the compression network or NDE), nor do we add extra layers. We use this framework for simplicity, since only one network needs to be trained and all training can be done end-to-end. Extensions to e.g. neural likelihood estimation, and further variants, will be explored in future work. We do not envision these would require any significant modifications, but would need several training stages as in Jeffrey et al. (2021).
We performed an initial exploration and hyperparameter tuning of architectures for data compression and density estimation. All model architecture and hyperparameter tuning was performed on the high-fidelity-only task. This ensures that our results are not biased toward improving transfer learning performance. We were partly motivated by the fact that common neural network architectures used for neural summarisation, such as CNNs, are well-suited for transfer between datasets. This stems from the inductive bias of CNNs, which encourages the learning of generic, transferable features that are then composed over a hierarchy of scales (Girshick et al., 2014; Yosinski et al., 2014; Kornblith et al., 2019). We found that the tailored CNN architecture from Villaescusa-Navarro et al. (2022), which was optimised for the CAMELS dataset using the hyperparameter optimisation framework optuna (Akiba et al., 2019), outperformed various standard architectures, such as ResNet (He et al., 2016) and ConvNext (Liu et al., 2022), both when pre-trained from natural image data or randomly initialised. We therefore used the CNN architecture from Villaescusa-Navarro et al. (2022) as our neural compression backbone, changing only the dimension of the final output layer to instead serve as a latent embedding. We found that using the CNN to compress matter density maps to larger latent dimension sizes slightly improved performance, so set the latent dimension to $128$.
We used a rational-quadratic neural spline flow (RQ-NSF, Durkan et al., 2019) as the NDE head of the network, implemented using the sbi Python package (Tejero-Cantero et al., 2022). We found this NDE architecture gave significant improvements over alternative popular choices, such as Masked Autoregressive Flows (MAFs, Papamakarios et al., 2017). We also found that inserting batch normalisation (Ioffe & Szegedy, 2015; Santurkar et al., 2018) layers between spline flow blocks (similar to the CNN architecture) substantially improved performance, both in pre-training and fine-tuning. An overview of the architecture is presented in Fig.˜2.
@@
Once the architecture was selected, we performed end-to-end training of the network on the NPE objective in Eq.˜2. We made a range of modifications that improved performance. We utilised a weight decay of $0.01$, which regularises the neural network by adding a small penalty term to the network weight magnitudes (Krogh & Hertz, 1991; Loshchilov & Hutter, 2017). This was particularly important for the fine-tuning stage, where regularisation while training on very small datasets was greatly beneficial. We used a short learning rate (LR) warm-up period, which has been found to improve deep learning model training (He et al., 2016; Goyal et al., 2017; Vaswani et al., 2017) with a number of posited explanations (see e.g., Gotmare et al., 2019; Kalra & Barkeshli, 2024). Larger batch-sizes (we selected 64) also improved performance for all models. We found that using a cyclic learning rate scheduler (Smith, 2017) as in Villaescusa-Navarro et al. (2022) improved performance when training on large datasets, i.e. $\mathcal{O}(10^{4})$ maps. Baseline experiments and pre-training were performed with a LR of $2\times 10^{-4}$. Fine-tuning was performed with a LR of $1\times 10^{-5}$ and an exponential decay scheduler. All models were trained using the AdamW optimizer (Loshchilov & Hutter, 2017).
We repeat all training runs six times, changing only the initialisation of the network, the random selection of $N$ maps from the training dataset, and the (random) order in which the training data is passed to the network. The model with the lowest validation loss from each run is saved and used for evaluation. All models pre-trained on $N$-body simulations used the entire $N$-body training suite, and models with the lowest validation loss were saved for fine-tuning and evaluation.
@@ 2.3.1 Posterior Accuracy
Figure 3: Inference results on the IllustrisTNG LH suite for a 2D posterior over $\Omega_{\textrm{m}}$ and $\sigma_{8}$. We compare the performance of the two approaches: training with only high-fidelity maps (blue) against the transfer learning approach (green), which uses $12000$ NN-body maps for pre-training. An accurate benchmark model trained on the entire IllustrisTNG LH suite training set is shown by the dashed red line. The accurate model (red) uses a cyclic learning rate scheduler, which performs slightly better than an exponential scheduler for large datasets, as detailed in Section˜$2.$ Panel a) shows the mean test posterior probability (MTPP), panel b) shows the calibration error $\mathcal{C}$ of the modelled posterior, defined in Eq.˜4, and panel c) shows the FoM, all as a function of high-fidelity dataset size $N$. Panel c) should be interpreted with the proviso that the model must be well-calibrated before the FoM measures genuine constraining power. All results show the mean and standard error over six independent training runs (including independent pre-training runs on the $N$-body simulations).
@@ 3 Results
We run a range of experiments comparing high-fidelity-only models against the transfer learning approach. We use $N$ to denote the number of IllustrisTNG 2D dark matter density maps used during the training stage. Of the dataset size $N$, $90\%$ is used as training data, while $10\%$ is used for validation data. All results report the mean performance on the test set.
@@ 3.1 CAMELS Multifield Dataset: LH suite
We compare the performance of each method over a range of IllustrisTNG dataset sizes. We present the results of the LH experiments in Fig.˜3. We find that small IllustrisTNG dataset sizes lead to poor performance when training from random network initialisation, whereas pre-training on $N$-body maps leads to good performance with very few simulations. For instance, a pre-trained model that is then fine-tuned with $N=200$ IllustrisTNG maps has higher MTPP than the high-fidelity-only approach with $N=6400$ maps.
@@
Insets: the overcoverage values per credibility level, see Eq.˜4 and the surrounding text for details. The shaded orange region highlights the discrepancy between the ideal and observed distribution of credibility levels, which is quantified by the calibration error metric introduced in Eq.˜4. Panel a) shows training from high-fidelity-only simulations, where models with small dataset sizes display significant overconfidence, and better calibration (though now mildly underconfident) is achieved at $N=6400$. Panel b) shows that transfer learning requires around $N=800$ IllustrisTNG maps to achieve good calibration.
@@
One complicating factor is the posterior calibration, quantified in Fig.˜3b. This demonstrates that despite the high test posterior probability, models fine-tuned with very few IllustrisTNG maps appear to be poorly calibrated. We find that acceptable calibration is achieved after $N=800$ fine-tuning maps, while training with only high-fidelity maps requires $N=6400$ for similar posterior calibration. We therefore find at least a factor of $8$ reduction in the number of simulations required to produce a performant, well-calibrated model of the posterior.
We present calibration curves from a range of dataset sizes across the two approaches in Fig.˜4. These show the standard cumulative distribution of observed credibility levels in the main panels, as well as the overcoverage distribution $\hat{p}_{i}/p_{i}$ in the insets. These reaffirm the calibration issues identified in Fig.˜3: training from scratch with fewer than $N=6400$ maps leads to significantly overconfident posteriors. The high overcoverage at $\{0,1\}$ (paired with the below-ideal coverage in the middle of the distribution) is a clear indicator of overconfidence, since it indicates that the true parameters occur at extreme credibility levels too often. On the other hand, the transfer learning models display a more minor form of bias and overconfidence until reaching around $N=800$ maps.
The FoM performance as a function of dataset size is shown in Fig.˜3c. When training using only high-fidelity simulations, low dataset sizes lead to low FoMs. This observation indicates that the features (or summary statistics) extracted by the CNN are not particularly informative. Additionally, the overconfident posteriors indicate limitations in the performance of the NDE, since it is incapable of producing trustworthy posteriors. We therefore conclude that both the CNN neural compressor and the density estimation model perform poorly with small training datasets.
@@
Three representative examples of inference on test cosmology maps are shown in Fig.˜5. We compare models produced using only high-fidelity maps ($N=800$), transfer learning ($N=800$), and an “accurate” reference model trained on the full IllustrisTNG training set. These examples are qualitatively consistent with the analysis presented above. The high-fidelity-only $N=800$ model gives very uninformative constraints compared to the other two posteriors. On the other hand, the fine-tuned model appears well-calibrated, and only slightly less constraining than the model trained with $\times 16.25$ more high-fidelity maps. We present a similar comparison with a high-fidelity model trained on $N=3200$ in Appendix˜A, demonstrating that even for larger high-fidelity dataset sizes, pre-training yields significantly improved posteriors.
@@
Appendix˜B presents a range of further tests into the model performance. We found that the pre-trained models performed very poorly on high-fidelity maps when no fine-tuning was performed (corresponding to $N=0$). We explored the quality of the fine-tuned CNN compressor by freezing the CNN and re-training the NDE with larger dataset sizes. This presented more evidence that the limiting factor at very low fine-tuning dataset sizes was the NDE. We also showed that the small performance gap between the “accurate” model and the transfer learning models with larger dataset sizes was caused by slightly worse compression.
@@
Another consideration was the possibility that the paired aspect of the datasets was responsible for the significant performance gains from lower-fidelity pre-training. While we did make any explicit use of the simulation pairs, this could still have had an impact depending on the training dynamics of the network. We tested this in Appendix˜B and found strong evidence that the paired aspect of the source and target dataset has no impact on our results. We therefore conclude that our transfer learning approach does not depend on paired datasets. In principle, this means that large, pre-existing simulation suites could be used as multifidelity datasets without the need to pair initial conditions and cosmological parameters.
@@
We found that fine-tuning with the entire IllustrisTNG suite produced a slightly weaker model than the “accurate” benchmark model (i.e. the transfer learning curve in Fig.˜3a does not intersect with the “accurate” performance). We analyse the convergence properties of the approaches further in Appendix˜B. We show that the slightly lower transfer learning MTPP plateau is partially due to the low learning rate (LR) used during fine-tuning, and a higher LR nearly recovers the “accurate” performance. This is consistent with the intuitive notion that a higher LR allows the training procedure to escape from the slightly sub-optimal region of the weight-space that is reached during pre-training. These observations are perfectly compatible with the study of Sharma et al. (2024), who found no clear benefits of transfer learning when using a large high-fidelity dataset.
@@
We present the degree of agreement between the posterior sample means $\hat{\theta}$ and the true cosmologies $\theta$ in Fig.˜6, broken down for $\sigma_{8}$ and $\Omega_{\textrm{m}}$. These indicate pre-training yields very large improvements in the posterior for both parameters. $\Omega_{\textrm{m}}$ and $\sigma_{8}$ are inferred with similar precision relative to the baseline over the entire test set.